OpenStudy (anonymous):

help me please If a varies directly as b, and a = 28 when b = 7, find b when a = 5

OpenStudy (anonymous):

a = k*b 28 = k * 7 find k

OpenStudy (anonymous):

is 39.2

OpenStudy (anonymous):

@surdawi god i hated these things

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

If a varies directly as b, and a = 28 when b = 7, find b when a = 5. 4/5 5/4 39.2

OpenStudy (anonymous):

is c

OpenStudy (anonymous):

just find the the k value in here 28 = k * 7

OpenStudy (anonymous):

is 4/5

OpenStudy (anonymous):

right, you got it